OpenStudy (anonymous):

\[\sqrt{3+2\sqrt{2}} - \sqrt{3-2\sqrt{2}}\] How do you solve this without using a calculator? How the real problem looks like is under this.

OpenStudy (anonymous):

\[\sqrt{3+2\sqrt{2}} - \sqrt{3-2\sqrt{2}}\]

OpenStudy (anonymous):

multiply by the conjugate

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

amistre do you know how?

OpenStudy (amistre64):

all cause I typoed it in the beginning eh ... maybe

OpenStudy (amistre64):

sqrt(3+2sqrt(2)) - sqrt(3-2sqrt(2)) = N ; ^2 3+2sqrt(2) + 3-2sqrt(2) -2sqrt((3+2sqrt(2))(3-2sqrt(2))) = N^2 6 -2sqrt[(3+2sqrt(2))(3-2sqrt(2))] = N^2 3 +2sqrt(2) 3 -2sqrt(2) ----------- 9 +6x -6x -8 ---------- 9 -8 = 1 6 -2sqrt(1) = N^2 6 -2(1) = N^2 6 - 2 = N^2 = 4 N = +- 2 ; but in this case its gonna be 2

OpenStudy (anonymous):

did you delete the other one? and the only thing i dont understand is going from steps 1 to 2

OpenStudy (amistre64):

well, i started out by making it equal some number "N" and then tried to solve for N the problem is that I squared each side and that has a tendency to bring in extra answers that dont apply

OpenStudy (anonymous):

ohh okay. because i know the answer is two but i didnt know how to prove it

OpenStudy (amistre64):

we can sift out the +- part with a bit of logic: sqrt(4+2) - sqrt(4-2) is the same set up sqrt(6) - sqrt(2) sqrt(6) is larger than sqrt(2) so itll be a positive result

OpenStudy (anonymous):

ok thanks!! and do you know how to solve\[\sqrt{2}+\sqrt{6} / \sqrt{2+\sqrt{3}}\] and the square root after the divisor is actually under the the first part but i dont know how to show it on here

OpenStudy (amistre64):

\frac{top}{bottom} \[\frac{top}{bottom}\]

OpenStudy (amistre64):

id just equate it to some unknown number N again and see what I could do to persuade it thru algebra

OpenStudy (anonymous):

what does this say that you typed on the third line?? 6 -2sqrt[(3+2sqrt(2))(3-2sqrt(2))] = N^2

OpenStudy (anonymous):

where did you get 3 +2sqrt(2) 3 -2sqrt(2) ----------- 9 +6x -6x -8 ---------- 9 -8 = 1 from??